If a ball is thrown upwards at 20 ft/sec, the height h after t seconds is given by h = -16t^2 + 20t. 

How many seconds does it take for the ball to reach its maximum height, and what is this height?
Answer by Alan3354(69443)

If a ball is thrown upwards at 20 ft/sec, the height h after t seconds is given by h = -16t^2 + 20t. 

How many seconds does it take for the ball to reach its maximum height, and what is this height?
---------------
The max ht is the vertex of the parabola
The line of symmetry is t = -b/2a = -20/-32
t = 5/8 seconds to reach max ht
------------
h(5/8) = -16*(5/8)^2 + 20*5/8
= -400/64 + 100/8 = 400/8
Max ht = 50 feet